A conductivity cell having a cell constant of 0.5 cm –1 filled with 0.02 M solution of KCl at  298 K gave a resistance of 20.2 Ω. The water used for preparing the solution had a conductivity  of 7.1 × 10–6 S cm–1. Calculate the molar conductivity of 0.02 M KCl solution.


1
Expert's answer
2021-03-08T06:00:24-0500

lmbda = k*1000/N

K = l/R*a

k = 0.5/20.2 = 0.025ohm^-1^cm^-1


lmbda = 0.025*1000/0.02 = 1250